### v1

@@ III-B Modeling the Keyframe Stream
where $f(x,\lambda)$ is PDF function, $\lambda$ is the expected rate of arrival of a new keyframe. This modeling of the keyframe stream gives us insights for designing a new sampling method that emphasizes recent frames while considering the keyframe arrival rate.
III-C Sampling with Shifted Exponential Distribution
The sampling problem we deal with is determining which keyframes we should sample rays/pixels from for each training iteration, given the limited number of total iterations for online training. As discussed in Section I, the naive uniform sampling approach does not sufficiently sample the later frames resulting in a rendering quality drop. Thus, we need a sampling method that generates more samples from the more recently received frames during training.
Before presenting our method, we first define how we represent time for NeRF training. Since the sampling is performed once per training iteration, and these iterations roughly take the same amount of time, we use the number of iterations to denote elapsed time. For instance, the timestamp $T_{i}$ of keyframe $I_{i}$ corresponds to the $T_{i}$-th iteration when the NeRF training routine first encounters $I_{i}$.
We seek to define a function $f$ that maps each keyframe to a sampling weight. The earlier frames that are far behind the current iteration should have lower weights and perform a more uniform sampling, while the newer frames should have higher weights. To ensure this function remains unaffected by the continuous increase in training iterations, we use relative time intervals as input for $f$, instead of absolute timestamps. Specifically, the time interval $D_{i}$ for a keyframe with timestamp $T_{i}$ at iteration $S$ is calculated as $D_{i}=S-T_{i}$, ensuring that newer frames consistently have lower $D_{i}$. Thus, the function $f(D)$ should be decreasing over time.
@@ III-C Sampling with Shifted Exponential Distribution
However, simply using this function would severely downweight the earlier frames. Because of the exponential decrease, the earlier frames will have sampling weights very close to zero and almost all the samples will come from the recent frames. The lack of access to the earlier frames could cause the forgetting issue [35] in NeRF training which could further cause the overall rendering quality drop [40, 44].
@@
Fig. 4 demonstrates the sampling results with varied $\alpha$ and $\beta$ values. Our proposed sampling function can gradually decrease the sampling weight of a newly received frame as training proceeds. With the help of the $\beta$ term, the sampling for earlier frames is similar to uniform sampling, providing sufficient training samples to avoid the forgetting issue.
TABLE I: Quantitative comparison on Replica scenes.
Method
off-0
off-1
off-2
off-3
off-4
rm-0
rm-1
rm-2
Avg.
PSNR normal-$\uparrow$
offline 34.31 33.52 27.78 27.61 29.44 26.80 28.16 29.42 29.63
uniform
33.23
32.42
26.88
27.09
28.06
25.96
27.47
29.07
28.77
imap frame
33.40
32.77
27.05
27.23
28.71
26.21
27.71
28.88
28.99
exp
33.51
32.77
27.12
27.24
28.51
26.27
27.67
29.42
29.06
loss only
33.75
33.51
27.32
27.02
28.27
26.30
27.81
29.09
29.13
imap loss 33.89 33.54 27.84 27.65 29.17 26.54 27.90 29.64 29.52
exp loss 34.43 33.71 27.82 27.88 29.23 26.63 28.09 29.79 29.70
SSIM normal-$\uparrow$
offline 0.921 0.901 0.882 0.873 0.901 0.776 0.819
0.879
0.869
uniform
0.905
0.885
0.876
0.862
0.890
0.752
0.815
0.880
0.858
imap frame
0.908
0.889
0.875
0.865
0.895
0.758
0.816
0.877
0.860
exp
0.909
0.884
0.879
0.867
0.894
0.760
0.816 0.883
0.861
loss only 0.913 0.899 0.880
0.863
0.891
0.760
0.819
0.879
0.863
imap loss 0.913
0.897
0.880 0.869 0.896 0.766 0.816 0.884
0.865
exp loss 0.919 0.901 0.883 0.872 0.898 0.768 0.821 0.887 0.869
LPIPS normal-$\downarrow$
offline 0.176 0.212
0.268
0.238 0.260 0.323 0.301 0.227 0.251
uniform
0.220
0.253
0.277
0.259
0.289
0.383
0.316
0.248
0.281
imap frame
0.223
0.232
0.278
0.259
0.275
0.362
0.313
0.247
0.274
exp
0.206
0.240
0.267 0.251
0.274
0.360
0.318
0.235
0.269
loss only
0.204
0.211 0.261
0.270
0.286
0.369
0.308
0.233
0.268
imap loss 0.202
0.213
0.262
0.253
0.273 0.356 0.305
0.238
0.263
exp loss 0.187 0.211
0.268
0.240 0.268 0.359 0.307 0.228 0.259
III-D Loss-Guided Active Ray Sampling
Similar to prior works that actively sample more rays/pixels on the regions with higher training loss [40, 5], our proposed frame sampling method in Sec. III-C can also be incorporated with a loss-guided active ray sampling method to improve the performance of the trained model further. The loss-guided active ray sampling method typically involves tracking the spatial loss distribution for each frame during training iterations. To achieve this, we divide each image into an $M\times M$ grid, where each grid patch maintains a running average L1 photometric loss of sampled training rays or pixels within that patch.
To incorporate loss active ray sampling with our shifted exponential frame sampling, we split the sampling into two parts. Suppose we need to sample $B$ rays for one iteration, we first use the sampling weight function (Eq. 4) to calculate the portion of frames $p_{r}$ that should be sampled from the frames that are received within the last $K$ training iterations. We then uniformly sample $B_{r}=\lfloor p_{r}B\rfloor$ rays from those recent frames according to their sampling weight. For the remaining $B-B_{r}$ rays, we sample them based on the patch loss distribution across all the rest unsampled frames.
Note that this loss active sampling is not well suited for real-captured outdoor scenes that contain significant background noise and many moving objects. In such cases, outliers often contribute to higher loss values, but optimizing these regions does not necessarily improve the overall model quality. In fact, it can degrade model performance by utilizing valuable sampling resources inefficiently. Hence, we simply use uniform ray sampling for noisy outdoor scenes.
IV Experiments
IV-A Implementation Details
To deploy our online NeRF framework, we use a desktop machine with an RTX4090 GPU as the remote server for NeRF training and a Jetson Xavier NX as the local edge processor, meeting SLAM computation needs efficiently with low energy consumption. The local end and remote server are connected through a WiFi network. We choose NeRFacto [11], a NeRF model that integrates many recent advancements and performance optimizations, as the default model in the experiments. We enable NeRFacto’s pose refinement feature [45] to make it more robust to less accurate camera poses from real-time tracking. We leverage a customized ORB-SLAM2 [12] as the on-device SLAM for the local end.
It is challenging to fairly compare different online approaches due to the impact of the randomness of the data generated by the local end and transmitted across the network. We therefore implement a simulation module on the remote end, which replays the keyframe stream logged by real-time SLAM systems. This approach guarantees that the keyframe stream exposed to the NeRF training process remains consistent across different training strategies for an accurate and fair comparison. The offline NeRF training in our experiments also uses the same set of keyframes but with full access to all the keyframes throughout the training. For all the evaluated online NeRF training strategies, we ensure that at any specific training iteration, the same number of keyframes is presented across different strategies. This is achieved by replaying a keyframe-iteration log, which is tracked by running the naive online NeRF training without any computation interference on the remote end. The online NeRF training stops upon receipt of all keyframes, with only a few additional training iterations (shorter than the average keyframe time interval). This tracked number of training iterations is then applied across all the evaluated training strategies. We evaluate scenes with PSNR, SSIM, and LPIPS metrics on the uniformly sampled frames along the camera trajectory to quantitatively evaluate the rendering quality.
TABLE II: Quantitative comparison on TnT scenes.
Scene Barn Train Truck
Method PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
offline 25.30 0.721 0.290 23.54 0.725 0.232 24.78 0.791 0.168
uniform
23.23
0.627
0.418
20.41
0.554
0.451
23.27
0.723
0.260
imap frame 23.91 0.660 0.368 21.40 0.603 0.378 23.59 0.737 0.232
exp 24.22 0.677 0.345 21.81 0.624 0.352 23.74 0.739 0.236
IV-B Datasets
We evaluate our method using the Replica [46] and Tanks and Temples [47] datasets. Replica contains various synthetic small-scale indoor scenes, we use the same camera trajectory of around 1-minute length from iMAP [40] for our online NeRF learning. The tanks and temples (TnT) dataset has a collection of high-resolution real-captured video recordings (3-7 minutes) of various indoor and outdoor scenes. We pick a small subset of outdoor scenes to showcase the capability of our method for challenging unbounded outdoor scenes. We utilize ORB-SLAM2 with downscaled frames (640x360) and a reduced frame rate (10 FPS) for more stable tracking performance on the TnT video input. We use RGB video streams of scenes for all evaluated methods.
office2
office3
room1
@@ IV-B Datasets
offline
uniform
imap frame
exp
imap loss
exp loss
Figure 6: Visual comparison of synthesized views from Replica scenes.
Barn
Truck
@@
offline
uniform
imap frame
exp
Figure 7: Visual comparison of NeRF-synthesized TnT scenes.
IV-C Evaluation
We first include offline NeRF training (“offline”) and online NeRF training with naive uniform sampling (“uniform”) as two fundamental baselines for our comparison. We then compare the methods with the keyframe sampling strategies. We implement a keyframe sampling method similar to iMAP (“imap frame”) that always samples $1/5$ rays of training batch from the most recent keyframe and uniformly samples the rest of rays from remaining keyframes. For indoor scenes from Replica, we additionally compare methods that leverage the loss-guided active sampling method discussed in Sec. III-D. The first baseline is the pure loss active sampling (“loss only”) which utilizes patch-wise loss distribution to sample rays. We also combine “imap frame” with loss active sampling to create a stronger iMAP baseline (“imap loss”). We use “exp” to denote our shifted exponential sampling method, and “exp loss” to denote “exp” with loss active sampling. We use $\alpha=2$, $\beta=4$ for our sampling methods.
IV-D Result Comparison
@@
Replica. In Table I and Fig. 6, we present quantitative and qualitative results for the Replica dataset, respectively. These results show that our shifted exponential sampling method improves the quality of renderings across various scenes. Our frame sampling method (“exp”) consistently does better than the naive uniform sampling and iMAP’s frame sampling. Furthermore, when combined with the loss-guided active sampling method, our method “exp loss” even outperforms the offline baseline on some replica scenes. The visual results in Fig. 6 demonstrate the noticeable quality disparities among different online training methods.
Tanks and Temples. We evaluate three complex outdoor scenes with results shown in Table II and Fig. 7. Our frame sampling method still consistently outperforms the naive uniform sampling method and iMAP frame sampling method on most metrics.
We conclude from these results that our shifted exponential sampling can effectively improve the rendering and visualization quality for online NeRF training from a streaming edge device.
IV-E Evaluation with Gaussian Splatting
To further demonstrate the applicability and effectiveness of our sampling method, we apply our sampling method to 3D Gaussian Splatting (3DGS) [48], a recently introduced point-based 3D representation method, for online 3D scene reconstruction. We evaluate the online 3DGS training on two indoor scenes from TnT dataset (with COLMAP [13] poses for this experiment). In our setting, 3DGS is initialized with 10K random points. The results in Table III indicate that our frame sampling method continues to enhance rendering quality, even when working with a substantially different 3D model. Conversely, iMAP’s frame sampling approach performs even worse than the basic uniform sampling method.
TABLE III: Metrics for online 3DGS training on TnT scenes.
Scene Church Meetingroom
Method PSNR SSIM LPIPS PSNR SSIM LPIPS
offline 23.33 0.812 0.251 26.88 0.902 0.162
uniform 22.49 0.790 0.276 26.05 0.884 0.187
imap frame
22.15
0.779
0.279
25.74
0.880
0.190
exp 22.55 0.788 0.273 26.24 0.888 0.183
V Conclusion
@@
We introduce a distributed framework, DISORF, to enable online 3D reconstruction and visualization of scenes that are captured from cameras on resource-constrained mobile robots. DISORF leverages the posed keyframes computed by on-device SLAM and higher computational capabilities by leveraging a remote server over the network to enable online NeRF training. We observe and address a key challenge with naive online NeRF training compared to the offline training is the unbalanced frame sampling during training that leads to significant loss in quality. The experimental results on various datasets and models demonstrate the effectiveness and applicability of our proposed sampling method. We believe our framework has the potential to enable more use cases and downstream real-time robotics tasks by leveraging the high-quality online visualization and 3D representation.

### v2

@@ V-A Modeling the Keyframe Stream
where $f(x,\lambda)$ is PDF function, $\lambda$ is the expected rate of arrival of a new keyframe. This modelling of the keyframe stream gives us insights for designing a new sampling method that emphasizes recent frames while considering the keyframe arrival rate.
V-B Sampling with Shifted Exponential Distribution
The sampling problem we address involves determining which keyframes to sample from for each training iteration, given the limited number of total iterations available for online training. As discussed in Section I, the naive uniform sampling approach does not sufficiently sample the later frames resulting in a rendering quality drop. Therefore, we need a sampling method that generates more samples from the more recently received frames during training.
Before presenting our method, we define how we represent time units for NeRF/3DGS training. Since sampling is performed once per training iteration, and these iterations take roughly the same amount of time, we use the number of iterations to denote elapsed time. For instance, the timestamp $T_{i}$ of keyframe $I_{i}$ corresponds to the $T_{i}$-th iteration when the training routine first encounters $I_{i}$.
We seek to define a function $f$ that maps each keyframe to a sampling weight. The earlier frames far behind the current iteration should have lower weights and be sampled more uniformly, while the newer frames should have higher weights. To ensure this function remains unaffected by the continuous increase in training iterations, we use relative time intervals as input for $f$, instead of absolute timestamps. Specifically, the time interval $D_{i}$ for a keyframe with timestamp $T_{i}$ at iteration $S$ is calculated as $D_{i}=S-T_{i}$, ensuring that newer frames consistently have lower $D_{i}$. Thus, the function $f(D)$ should be descending over time.
@@ V-B Sampling with Shifted Exponential Distribution
However, simply using this function would severely downweight the earlier frames. Because of the exponential decrease, the earlier frames will have sampling weights very close to zero and almost all the samples will come from the recent frames. The lack of access to the earlier frames could cause the forgetting issue [38] in the training which could further cause the overall rendering quality drop [27, 39].
@@
Fig. 5 demonstrates the sampling results with varied $\alpha$ and $\beta$ values. Our proposed sampling function can gradually decrease the sampling weight of a newly received frame as training proceeds. With the help of the $\beta$ term, the sampling for earlier frames is similar to uniform sampling, providing sufficient training samples to avoid the forgetting issue. We empirically find a default setting of $\alpha=2,\beta=4$ would give a decent sampling balance for the online training.
VI Experiments
VI-A Setup
To deploy our online reconstruction framework, we use a desktop machine with an RTX4090 GPU as the remote server for reconstruction training. We use a Jetson Orin Nano board as our local edge processor, meeting the SLAM computation needs with low energy consumption.
We deploy a customized ORB-SLAM2 [9] as our on-device SLAM module for the local end. As mentioned in Sec. IV, we evaluate our online training with both NeRF and 3DGS as the 3D representation model on the remote server. We also enable the differentiable pose refinement [40] to make our training more robust to less accurate camera poses from real-time tracking.
It is challenging to fairly compare different online approaches due to the impact of the randomness of the data generated by the local end and transmitted over the network. Therefore, we implement a simulation module on the remote end, replaying the keyframe stream logged by real-time SLAM systems. This approach guarantees that the keyframe stream exposed to the training process remains consistent across different training strategies for an accurate and fair comparison. For all the evaluated online training strategies, we ensure that the same number of keyframes is presented across different strategies at any specific training iteration. This is achieved by replaying a keyframe-iteration log, which is recorded by running the naïve online training without any computation interference on the remote end. The online training stops upon receipt of all keyframes, with only a few additional training iterations (less than 10 seconds). The recorded number of training iterations is then applied across all the evaluated training strategies. The offline model training in our experiments uses the same set of keyframes but with full access to all the keyframes at the beginning of training. To quantitatively evaluate the rendering quality, we evaluate scenes with PSNR, SSIM, and LPIPS metrics on the uniformly sampled frames along the camera trajectory.
TABLE I: Quantitative results on Replica dataset, scores are averaged over 8 scans.
Method NeRF Method 3DGS
PSNR SSIM LPIPS PSNR SSIM LPIPS
offline 29.63 0.869 0.251 offline 31.03 0.907 0.173
uniform 28.77 0.858 0.281 uniform 29.39 0.877 0.200
imap 28.99 0.860 0.274 imap 29.07 0.875 0.205
ours 29.06 0.861 0.269 ours 29.89 0.882 0.190
imap+loss 29.52 0.865 0.263 imap+mvs 29.78 0.883 0.187
ours+loss 29.70 0.869 0.259 ours+mvs 30.01 0.883 0.185
VI-B Datasets
We evaluate our method using the Replica [41] and Tanks and Temples [42] datasets. Replica contains various synthetic small-scale indoor scenes. We use the same camera trajectory of around 1-minute length from iMAP [27] for our online NeRF learning. The tanks and temples (TnT) dataset has a collection of high-resolution real-captured video recordings (3-7 minutes) of various indoor and outdoor scenes. We pick a small subset of outdoor scenes to showcase the capability of our method for challenging unbounded outdoor scenes. We utilize ORB-SLAM2 with downscaled frames (640x360) and a reduced frame rate (10 FPS) for more stable tracking performance on the TnT video input. We use RGB video streams of scenes for all evaluated methods. We also captured 2 scenes with a handheld compute board, the results are shown in the supplement video [^3] https://www.youtube.com/watch?v=qphP5QnMp6w.
TABLE II: Quantitative results on Tanks&Temples scenes.
Method Barn Train Truck
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
NeRF offline 25.30 0.721 0.290 23.54 0.725 0.232 24.78 0.791 0.168
uniform 23.23 0.627 0.418 20.41 0.554 0.451 23.27 0.723 0.260
imap 23.91 0.660 0.368 21.40 0.603 0.378 23.59 0.737 0.232
ours 24.22 0.677 0.345 21.81 0.624 0.352 23.74 0.739 0.236
3DGS offline 26.58 0.852 0.142 21.70 0.792 0.178 24.74 0.871 0.109
uniform 24.22 0.782 0.172 20.46 0.703 0.228 22.26 0.783 0.153
imap 24.48 0.798 0.160 20.44 0.701 0.239 22.47 0.791 0.152
ours 24.67 0.803 0.162 20.73 0.722 0.217 22.75 0.803 0.143
@@ VI-B Datasets
NeRF office2 [Uncaptioned image]
office3 [Uncaptioned image]
@@
[…]
3DGS office0 [Uncaptioned image]
room0 [Uncaptioned image]
[…]
Figure 6: Visual comparisons on Replica scenes.
[…]
NeRF Barn [Uncaptioned image]
Truck [Uncaptioned image]
[…]
3DGS Barn [Uncaptioned image]
Train [Uncaptioned image]
[…]
Figure 7: Visual comparisons on Tanks&Temples scenes.
VI-C Evaluation
We first include offline training (“offline”) and online training with naïve uniform sampling (“uniform”) as two fundamental baselines for our comparison. We then include the methods with different frame sampling strategies. We implement a frame sampling method similar to iMAP [27] (“imap frame”) for both NeRF and 3DGS. iMAP always samples $20\%$ rays of the NeRF training batch from the most recent keyframe and uniformly samples the rest of the rays from the remaining keyframes. Since 3DGS training does not involve ray sampling, we instead sample the most recent keyframe with the probability of $20\%$, and $80\%$ for the remaining keyframes. We use “ours” to denote our shifted exponential sampling method. We also combine our proposed frame sampling with other training enhancement methods, to demonstrate its versatility and generalizability. For NeRF-based online reconstruction, we combine our frame sampling with the loss-guided ray sampling [27, 6]. This method tracks loss distribution by dividing each image into small patches and maintaining a running average photometric loss for each patch. Rays are then weighted-sampled based on tracked loss. We use “+loss” to denote methods with loss-guided ray sampling. For 3DGS-based online reconstruction, we also test our frame sampling with enhanced geometry prior [43]. We leverage DUSt3R [44], an efficient state-of-the-art multi-view stereo (MVS) method, to add and initialize Gaussian as estimated geometric priors of the 3D targets. We run DUSt3R on 4 newly received adjacent keyframes ($\sim$3 sec per DUSt3R inference) once every 500 iterations during the first 3000 training iterations to obtain MVS points to be used for adding new Gaussian points. We use “+mvs” to denote methods enhanced with MVS points.
VI-D Result Comparison
@@
Replica. In Table I and Fig. 6, we present quantitative and qualitative results for Replica dataset. These results demonstrate that our shifted exponential sampling method improves the quality of renderings across various scenes. Our frame sampling method (“ours”) consistently performs better than the naïve uniform sampling and iMAP’s frame sampling. Furthermore, when combined with other training enhancement methods (e.g., “+loss” and “+mvs”), our frame sampling can further improve the resulting quality. Our method “ours+loss” even outperforms the offline NeRF baseline on some scenes. The visual results in Fig. 6 highlight the noticeable quality differences among the different online training methods. Comparing NeRF and 3DGS, the 3DGS-based methods achieve higher rendering quality. Due to 3DGS’s efficient rasterization-based rendering, it can run up to 3 times more training iterations compared to NeRF training in the same timeframe. This efficiency allows 3DGS to reconstruct sharper images with more fine-grained details, particularly for the relatively simple Replica indoor scenes.
Tanks and Temples. We evaluate three complex outdoor scenes with results shown in Table II and Fig. 7. Our frame sampling method still consistently outperforms the naïve uniform sampling method and iMAP frame sampling method on most metrics. Unlike the results on the Replica dataset, NeRF methods achieve similar or even better quantitative results than 3DGS on these unbounded outdoor scenes. This is because the NeRF method, assisted by spatial contraction as an inductive bias, can better reconstruct 360 unbounded scenes. Due to the lack of reliable geometry priors, reconstructing distant objects in unbounded scenes, especially background objects, is challenging for 3DGS. Therefore, the NeRF-based method is more advantageous for robust reconstruction in our online setup.
VII Conclusion
@@
We introduce DISORF, a distributed framework that enables online 3D reconstruction and visualization of scenes captured by cameras on resource-constrained mobile robots. DISORF leverages posed keyframes computed by on-device SLAM and the computational power of a remote server to facilitate online NeRF/3DGS training. One key challenge we observed with naïve online training compared to offline training is the unbalanced frame sampling during training, leading to a significant loss in quality. To address this challenge, we propose a shifted exponential frame sampling method that adaptively gives more weight to more recent frames while still sufficiently sampling earlier frames. Experimental results on various datasets and models demonstrate the effectiveness and applicability of our proposed sampling method. We believe our modular framework can enable more real-time use cases and downstream robotics tasks by leveraging high-quality online visualization and 3D representation.